OpenStudy (anonymous):

In a collection of dimes and nickels, there are 36 coins. The total value of the coins is $2.90. How many coins of each type are there?

OpenStudy (anonymous):

You set up the problem like... D+N=36 10D+5N=2.90

OpenStudy (anonymous):

Where D is dimes and N is nickels

OpenStudy (anonymous):

thanks for the help!! 22 + 14= 36 0.10(22) + 0.05(14) 2.20 + 0.70 = $2.90

OpenStudy (anonymous):

Sure Sure!